OpenStudy (anonymous):

Write a rule for the linear function with slope m and y- intercept b. m = 2, b=- 4

ganeshie8 (ganeshie8):

slope = \(m\) yintercept = \(b\) the line equation is : \(\large y = mx + b\)

ganeshie8 (ganeshie8):

simply plugin the given numbers

OpenStudy (anonymous):

can you explain please

ganeshie8 (ganeshie8):

plug in the given values : \(\large m = 2\), and \(\large b=- 4\), in the above line equation : \(\large y = mx + b\)

ganeshie8 (ganeshie8):

you would get : \(\large y = 2x - 4\)

ganeshie8 (ganeshie8):

^^thats the rule for the linear function

OpenStudy (anonymous):

have you ever heard the term y=mx+c

OpenStudy (anonymous):

no